# One Perturbation is Enough: On Generating Universal Adversarial Perturbations against Vision-Language Pre-training Models

### v2

@@ Appendix D Multimodal Alignment Destruction
To provide more intuitive evidence that our C-PGC successfully destroys the image-text alignment relationship, we compute the distance between the encoded image and text embeddings before and after applying the UAP. Concretely, for an input pair (v, t), we calculate the relative distance by:
@@
| Ours | 6.41 | 12.15 | 13.64 | 4.35 | 1.71 | 1.96 |
Appendix E More Experimental Results
Visual entailment tasks. Given an image and a textual description, visual entailment involves determining whether the textual description can be inferred from the semantic information of the image. We align with previous VLP attacks (Zhang et al., 2022; Wang et al., 2024) and conduct experiments on the SNLI-VE (Xie et al., 2019) dataset using the ALBEF and TCL models. Note that the Baseline represents the clean performance of the target model on the clean data and the orange and green indicate ALBEF and TCL as source models respectively. The results presented in Figure 7 reveal that C-PGC obtains impressive attack effects by decreasing the average accuracy by nearly 20%. Notably, (Do et al., 2020) has reported a large number of annotation errors in the labels of the SNLI-VE corpus used for VE tasks. Therefore, the presented results are only for experimental integrity and reference purposes.
| Setting | Source | ALBEF | TCL | X-VLM | CLIP | CLIP | BLIP | ||||||
| TR | IR | TR | IR | TR | IR | TR | IR | TR | IR | TR | IR | ||
| CLIP | 34.78 | 50.42 | 37.17 | 51.24 | 36.81 | 50.87 | 63.07 | 70.92 | 92.48 | 93.66 | 41.81 | 55.12 | |
| CLIP | 14.29 | 36.57 | 20.5 | 41.7 | 15.55 | 37.06 | 41.63 | 62.87 | 86.53 | 88.73 | 17.98 | 44.31 | |
Cross-domain scenarios. We proceed to discuss the attack performance of the proposed algorithm in a more challenging scenario where there is an obvious distribution shift between the training dataset and the test samples. Specifically, we generate universal adversarial perturbations based on MSCOCO or Flickr30K and evaluate them accordingly on the other dataset. We present the attack success rates on the retrieval tasks across six models in Table 7. It can be observed that the domain gap indeed brings a negative effect on the attack performance. However, our method still maintains excellent ASR in most cases, unveiling its outstanding cross-domain transferability.
Results of R@5 and R@10. As aforementioned, we supplement the ASR of the ITR tasks based on R@5 and R@10 metrics and provide the attack success rates in Table 8. Obviously, our proposed C-PGC still consistently attains better performance than the baseline method GAP, regardless of the evaluation measurements for retrieval results.
| Dataset | Source | Method | ALBEF | TCL | X-VLM | CLIP | CLIP | BLIP | ||||||
| TR | IR | TR | IR | TR | IR | TR | IR | TR | IR | TR | IR | |||
| Flickr30K (R@5) | ALBEF | GAP | 55.71 | 73.86 | 8.01 | 10.54 | 1.2 | 4.84 | 4.46 | 15.24 | 8.28 | 20.27 | 5.33 | 10.77 |
| TCL | GAP | 17.64 | 20.09 | 77.89 | 74.53 | 0.9 | 4.2 | 4.25 | 15.48 | 8.6 | 20.25 | 8.65 | 13.16 | |
| X-VLM | GAP | 6.21 | 7.45 | 4.9 | 7.96 | 81.6 | 77.23 | 6.11 | 18.33 | 17.41 | 28.35 | 5.03 | 8.61 | |
| CLIP | GAP | 2.81 | 6.86 | 4.2 | 8 | 1.7 | 6.1 | 75.64 | 82.56 | 24.2 | 37.68 | 4.33 | 9.97 | |
| CLIP | GAP | 2.81 | 7.41 | 5.81 | 9.27 | 2.4 | 7.01 | 9.33 | 19.64 | 57.63 | 69.33 | 4.02 | 9.76 | |
| Ours | 3.01 | 19.09 | 5.11 | 22.7 | 3.3 | 23.07 | 17.41 | 41.17 | 61.57 | 74.32 | 6.74 | 25.16 | ||
| BLIP | GAP | 3.41 | 7.01 | 3.7 | 7.45 | 1 | 4.4 | 4.46 | 14.43 | 6.79 | 18.67 | 39.13 | 68.02 | |
| MSCOCO (R@5) | ALBEF | GAP | 74.43 | 78.62 | 37.99 | 30.08 | 5.56 | 7.19 | 14.26 | 17.11 | 15.58 | 21.62 | 23.73 | 23.26 |
| TCL | GAP | 41.48 | 32.59 | 92.54 | 87.81 | 6.46 | 8.08 | 16.09 | 18.47 | 17.98 | 24.3 | 29.9 | 28.64 | |
| X-VLM | GAP | 12.29 | 11.64 | 13.43 | 10.99 | 90.8 | 83.05 | 20.02 | 23.4 | 37.72 | 40.09 | 12.64 | 12.04 | |
| CLIP | GAP | 18.78 | 17.46 | 20.38 | 17.37 | 16.81 | 15.15 | 95.21 | 93.04 | 62.77 | 62.62 | 18.48 | 19.48 | |
| CLIP | GAP | 13.54 | 14.09 | 14.42 | 14.14 | 11.02 | 12.03 | 25.27 | 24.98 | 88.67 | 88.83 | 12.8 | 14.98 | |
| Ours | 16.25 | 30.07 | 20.96 | 34.15 | 16.58 | 30.23 | 48.04 | 56.66 | 91.54 | 88.96 | 21.37 | 38.85 | ||
| BLIP | GAP | 23.62 | 24.22 | 22.96 | 18.43 | 9.93 | 9.75 | 19.2 | 22.24 | 24.99 | 30.19 | 62.75 | 66.88 | |
| Flickr30K (R@10) | ALBEF | GAP | 51.6 | 71.17 | 5.8 | 6.65 | 0.6 | 2.7 | 1.42 | 9.82 | 4.19 | 13.81 | 3.71 | 7.01 |
| TCL | GAP | 14.9 | 14.29 | 73.26 | 70.49 | 0.6 | 2.22 | 2.13 | 9.73 | 4.29 | 13.45 | 5.92 | 9.51 | |
| X-VLM | GAP | 4.1 | 4.81 | 2.7 | 4.51 | 76.5 | 72.58 | 3.65 | 11.63 | 10.84 | 18.91 | 3.01 | 5.29 | |
| CLIP | GAP | 2.1 | 4.04 | 2.6 | 4.81 | 1 | 3.79 | 63.29 | 77.83 | 17.89 | 28.11 | 2.31 | 6.12 | |
| CLIP | GAP | 2.3 | 4.23 | 3.3 | 5.56 | 1.4 | 4.41 | 4.56 | 12.35 | 49.86 | 62.17 | 2.21 | 6.63 | |
| Ours | 2.5 | 13.56 | 3.7 | 16.87 | 1.7 | 17.38 | 9.93 | 32.4 | 53.27 | 66.34 | 3.91 | 19.62 | ||
| BLIP | GAP | 2 | 3.98 | 1.5 | 4.17 | 0.2 | 2.24 | 1.93 | 8.38 | 3.68 | 12.48 | 36.81 | 67.22 | |
| MSCOCO (R@10) | ALBEF | GAP | 69.78 | 76.24 | 32.04 | 23.7 | 3.16 | 4.91 | 10.07 | 13.16 | 12.55 | 16.6 | 19.1 | 19.96 |
| TCL | GAP | 34.36 | 25.76 | 90.65 | 85.27 | 4.01 | 5.44 | 11.77 | 14.88 | 13.66 | 18.89 | 24.91 | 24.67 | |
| X-VLM | GAP | 7.66 | 7.65 | 8.07 | 7.27 | 88.3 | 79.85 | 15.63 | 18.77 | 31.79 | 33.54 | 8.24 | 9.37 | |
| CLIP | GAP | 13.13 | 12.39 | 14.05 | 12.13 | 10.68 | 10.89 | 93.72 | 91.51 | 57.39 | 56.47 | 13.35 | 15.64 | |
| CLIP | GAP | 9.02 | 10.08 | 9.06 | 9.89 | 6.97 | 8.25 | 18.78 | 20.11 | 87.6 | 83.92 | 8.53 | 11.91 | |
| Ours | 10.68 | 22.98 | 14.19 | 27.21 | 10.62 | 23.96 | 39.89 | 49.62 | 88.74 | 85.18 | 15.97 | 33.25 | ||
| BLIP | GAP | 17.76 | 18.98 | 16.32 | 13.13 | 6.21 | 6.44 | 13.97 | 17.39 | 19.73 | 24.37 | 57.99 | 65.49 | |
Ablation study of the text perturbation. We introduce another variant C-PGCt that cancels the perturbation from the text side to investigate the contribution of image perturbation, text perturbation, and their synergy. The comparison results of C-PGC, C-PGCt, and GAP using ALBEF as the surrogate model are shown in Table 9.
We find that when merely applying image perturbations (C-PGCt), our design still outperforms GAP with notable improvements, validating the proposed techniques’ effectiveness in enhancing the image perturbation. Moreover, the superiority of C-PGC over C-PGCt indicates that the incorporation of textual perturbations can further boost the universal attacks on the basis of C-PGCt since the text perturbation facilitates the deconstruction of the learned cross-modal alignment.
| Method | ALBEF | TCL | X-VLM | CLIP | CLIP | BLIP | ||||||
| TR | IR | TR | IR | TR | IR | TR | IR | TR | IR | TR | IR | |
| GAP | 69.78 | 81.59 | 22.15 | 29.97 | 6.61 | 18.37 | 23.4 | 37.54 | 29.92 | 44.29 | 16.09 | 28.12 |
| C-PGCt | 86.74 | 86.3 | 50.1 | 50.2 | 10.87 | 21.53 | 26.28 | 39.3 | 33.42 | 48.32 | 31.55 | 36.77 |
| C-PGC | 90.13 | 88.82 | 62.11 | 64.48 | 20.53 | 39.38 | 43.1 | 65.93 | 54.4 | 72.51 | 44.79 | 56.36 |
Appendix F Limitations and Future Directions
@@
Overlook of interactions between perturbations and . The proposed framework generates universal perturbations for image and text respectively based on the designed multimodal and unimodal losses. Despite the remarkable attack performance, it does not consider the interactions and synergy between the perturbations and during optimization, which have been leveraged in several previous attacks (Lu et al., 2023; Wang et al., 2024) to improve performance. In future research, this limitation can be explored as a potential mechanism to further strengthen the attacks.
Textual Semantic Inconsistency. To ensure the stealthiness of textual attacks, we set the perturbation budget , i.e., only one word is modified. Nonetheless, this perturbation is still relatively likely to be detected since it is considerably difficult to maintain semantic consistency between the clean and the perturbed sentences when applying only one universal word to hundreds of various sentences. For instance, when the adversarial word is getting (verb) and the word to replace is bed (noun), the resultant adversarial text becomes A getting is unmade with pillows on it, which is somehow strange and likely to attract the attention of defense systems or humans. This is an inevitable common issue for the whole research area of textual UAP. A possible solution is to add perturbations to the word embeddings instead of the input sentence. However, this only applies to white-box scenarios given that attackers have no access to the word embeddings in the black-box settings.
Appendix G More Visualization Results
@@
This section presents a rich visual analysis of the proposed attack on a series of downstream tasks. Specifically, we generate the UAP and conduct attacks on the ALBEF model in the visual grounding (VG) task. As illustrated in Figure 8, the prediction bounding boxes exhibit a notable deviation from the clean predictions, verifying that our generated adversarial samples significantly interfere with the multimodal alignment. In the visual entailment (VE) task, we employ BLIP as the victim model and present the results in Figure 9. These qualitative visulizations again demonstrate the remarkable attack effects of our proposed method on various downstream tasks.

### v3

@@ Appendix F Multimodal Alignment Destruction
To provide more intuitive evidence that our C-PGC successfully destroys the image-text alignment relationship, we compute the distance between the encoded image and text embeddings before and after applying the UAP. For an input pair (v, t), we calculate the relative distance by:
@@
| ETU | 5.26 | 6.03 | 13.14 | 2.90 | 0.25 | 1.14 | |
| C-PGC | 6.41 | 12.15 | 13.64 | 4.35 | 1.71 | 1.96 |
Appendix G Discussions and Future Directions
@@
Overlook of interactions between perturbations and . The proposed framework generates universal perturbations for image and text respectively based on the designed loss functions. Despite the remarkable performance, it does not consider interactions between and during the optimization. Future research can explore this limitation as a potential mechanism to further strengthen the attack.
Textual Semantic consistency. To ensure the stealthiness of text attacks, we set the perturbation budget , i.e., only one word is modified. Although C-PGC obtains superior semantic similarity to previous sample-specific methods, there is still room to improve from the proposed C-PGC. Moreover, future studies can consider similarity preservation strategies by applying more effective constraints during the generator training or adversarial sentences post-processing to facilitate a more stealthy attack.
Appendix H More Visualization Results
@@
This section presents a rich visual analysis of the proposed attack on a series of downstream tasks. Specifically, we first provide the visualization of the image retrieval task using the MSCOCO dataset in Fig. 7. Besides, we generate the UAP using the ALBEF model for the visual grounding (VG) task. As illustrated in Fig. 8, the prediction bounding boxes exhibit a notable deviation from the clean predictions, verifying that our generated adversarial samples significantly interfere with the multimodal alignment. In the visual entailment (VE) task, we employ BLIP as the victim model and present the results in Fig. 9. These qualitative visualizations again demonstrate the remarkable attack effects of our proposed method on various downstream tasks.